# Second-Order Fine-Tuning without Pain for LLMs: A Hessian Informed Zeroth-Order Optimizer

### v1

@@ 5.3 Training with Non-Differentiable Objectives
Our proposed HiZOO employs gradient estimation to update parameters, allowing for the use of non-differentiable objectives for training. Following the setting of MeZO (Malladi et al., 2023), we conducted extensive experiments using F1 as objectives. The results presented in Table 4 indicate that our method outperforms MeZO by 6.36% on F1 on average. It’s worth noting that on TREC the improvement is even more significant (over 20%) and the original inability to converge using MeZO is solved.
@@
Table 4: HiZOO with non-differentiable objectives. For classification ($k=512$), we use HiZOO with full-parameter and optimize accuracy; for SQuAD (1,000 examples), we use HiZOO (prefix) and F1.
5.4 Memory Usage and Training Time Analysis
Figure 5: Experiments on influence of the value of Smooth scale $\alpha_{t}$ in Eq. (4). Here we use HiZOO (prefix) to fine-tune Roberta-large on SNLI dataset ($K=512$). More explored experiments can be found in Appendix G.2.
As the storage of the Hessian information, HiZOO increases a little memory usage compared to MeZO, but it still holds great advantages over FT or FT (prefix) (refer to Appendix F for detailed numbers). Specifically, using the same GPUs, HiZOO allows for tuning a model that is 5 times larger than what is feasible with FT on average, as shown in Table 5.
We also compare the training time efficiency between our HiZOO and MeZO. As we know, HiZOO requires 3 forward passes while MeZO requires 2 forward passes. Take fine-tuning RoBERTa-large as an example, although HiZOO requires $3/2$ the time of MeZO per step, HiZOO requires $1/5$ the training steps of MeZO. As a result, HiZOO can significantly reduce the overall training time by 70%. What’s more, HiZOO may converge with the smaller loss than MeZO and achieve a higher accuracy in many cases.
Hardware Largest OPT that can fit
FT FT-prefix MeZO HiZOO
1$\times$×A100 (80GB) 2.7B 6.7B 30B 13B
2$\times$×A100 (160GB) 6.7B 13B 66B 30B
4$\times$×A100 (320GB) 13B 30B 66B 66B
8$\times$×A100 (640GB) 30B 66B 175B 175B
Table 5: Largest OPT models that one can tune with specific hardwares and algorithms.
5.5 Influence of Smooth Scale $\alpha_{t}$
@@
To assess the robustness of the optimizer, a grid search is conducted to evaluate the sensitivity of hyper-parameter $\alpha_{t}$ on RoBERTa-large (350M). Figure 5 illustrates that the smoothing scale $\alpha_{t}$ is instrumental in controlling the convergence rate during the training. As $\alpha_{t}$ is incrementally increased from zero, the training loss decreases faster. However, excessively large values of $\alpha$ may impede convergence or even make training fail due to overflow or gradient explosion. The experimental results indicate that $\alpha=1e-6$ yields better performance in most cases.
5.6 Trajectory Visualization on Test Functions
In this section, we evaluate the performance of the Adam, MeZO, and our HiZOO on 3 test functions with heterogeneous curvatures across different parameters:
  • Function (a)  [^1] Figure 1 is plotted from Function (a) (Liu & Li, 2023).: $f(x,y)=8(x-1)^{2}(1.3x^{2}+2x+1)+0.5(y-4)^{2}$
  • Function (b): $f(x,y)=\lvert x\rvert+\lvert y\rvert$
  • Function (c): $f(x,y)=10000x^{2}+y^{2}$
The 2D trajectories in training are illustrated in Figure 6. HiZOO and Adam are both capable of achieving convergence on Functions (a), (b), and (c), yet HiZOO consistently converges more quickly than Adam. MeZO faces challenges, only achieving effective convergence in either the $x$ or $y$ dimension, but not both, indicating a limitation in handling functions with heterogeneous curvatures. Particularly, Function (c) with its high condition number, showcases HiZOO’s remarkable ability to quickly navigate towards convergence, outperforming other methods in efficiently dealing with heterogeneous curvatures between parameter dimensions.
@@ 5.6 Trajectory Visualization on Test Functions
Figure 7: Loss curves on Function (a) with different estimation times $n$ per step. Trajectory visualization can be found in Appendix G.1.
5.7 Influence of Estimation Times $n$ Per Step
Here we conduct experiments on Function (c) to explore the influence of estimation times $n$ per step. As shown in Figure 7, the bigger $n$ will decrease the variance of the diagonal Hessian estimation per step and reduces the overall training steps, but it will cause much more computation per step at the same time.
@@ 6 Conclusion
In this work, we introduced HiZOO optimizer, a diagonal Hessian enhanced zeroth-order optimizer for fine-tuning large language models across many tasks and scales. To our knowledge, HiZOO is the first ZO-SGD optimizer that incorporates diagonal Hessian for fine-tuning LLMs. Further experiment shows that HiZOO converges in much fewer steps than MeZO and achieves better performance, including LoRA, prefix and full-parameter tuning. We also conduct visualization experiments on popular test functions to show that HiZOO performs better than MeZO and comparably with Adam.
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.1 First-order Optimizer Used in LLMs
Optimization methods have consistently been a popular research domain, encompassing techniques such as Gradient Descent (GD), Momentum, Adagrad (Duchi et al., 2011), ADADELTA (Zeiler, 2012), and Newton’s method, which have been instrumental in advancing fields like computer vision. However, the emergence of large-scale models, characterized by their massive parameter counts and intricate architectures, has challenged the efficacy of conventional optimization methods for training tasks. Amidst this landscape, Adam (Kingma & Ba, 2015) has emerged as the preferred choice for its ability to rapidly converge, making it particularly suitable for the training and fine-tuning large models. Then AdamW (Loshchilov & Hutter, 2019) was proposed to add a weight decay coefficient to alleviate over-fitting. Notwithstanding these advancements, a limitation persists with these optimizers: they have an implicit batch size ceiling. Exceeding this threshold can provoke extreme gradient updates, thus impeding the convergence rate of the models. This bottleneck is particularly problematic in the context of large-model training, which typically necessitates substantial batch sizes. To circumvent this constraint, LAMB (You et al., 2020) was devised to apply principled layer-wise adaptation strategy to accelerate the training of large models employing large batches.
@@ A.2 Hessian Based First-Order Optimizer
Compared with first-order optimizers, second-order optimizer considers second-order information in the process of gradient calculation. As a result, it has more abundant information to guide gradient descent and is considered to be more promising. Previous studies utilized curvature information to pre-condition the gradient (BROYDEN, 1970; Nesterov & Polyak, 2006; Conn et al., 2000). Subsequently, Magoulas et al. (1999) applied diagonal Hessian as the pre-conditioner, which greatly promotes the landing of second-order optimizer in the field of deep learning. Martens (2010) approximated the Hessian with conjugate gradient. Schaul et al. (2013) utilized diagonal Hessian to automatically adjust the learning rate of SGD during training. Pascanu & Bengio (2014) extended natural gradient descent to incorporate second order information alongside the manifold information and used a truncated Newton approach for inverting the metric matrix instead of using a diagonal approximation of it. EVA  (Zhang et al., 2023) proposed to use the Kronecker factorization of two small vectors to approximated the Hessian, which significantly reduces memory consumption. AdaHessian (Yao et al., 2021) incorporates an approximate Hessian diagonal, with spatial averaging and momentum to precondition the gradient vector.
Although great progress has been made in the research of second-order optimizer, it has not been widely used because of the extra computation and memory cost when gradient updating, and this situation is extremely serious in the training of large language models. Based on the above dilemma, Anil et al. (2021) proposed to offload Hessian computation to CPUs, and George et al. (2018) utilized ResNets and very large batch size to approximate the Fisher information matrix. Sophia (Liu & Li, 2023) was the first to apply second-order optimizer and achieve a speed-up on large language models in total compute successfully.
@@ A.3 Zeroth-Order Optimizer
Zeroth-order optimization, is also known as derivative-free or black-box optimization. There have been many one-point gradient estimators in past works (FairScale authors, 2021; Spall, 1997; Vakhitov et al., 2009; Spall, 1992; Jamieson et al., 2012; Agarwal et al., 2009; Raginsky & Rakhlin, 2011; Wang et al., 2020). However, cursory experiments with one such promising estimator (Zhang et al., 2022b) reveal that SPSA outperforms other methods.
In previous works, it appears in a wide range of applications where either the objective function is implicit or its gradient is impossible or too expensive to compute. For example, Tang et al. (2021) and Hajinezhad & Zavlanos (2018) consider derivative-free distributed algorithms for non-convex multi-agent optimization. ZO-BCD(Cai et al., 2021), ZOO(Chen et al., 2017), ZO-signSGD (Liu et al., 2019a) and ZO-HessAware (Ye et al., 2019) utilize zeroth-order stochastic optimization to generate black-box adversarial example in deep learning.
Beyond that, MeZO (Malladi et al., 2023) firstly adapted the classical ZO-SGD method to fine-tune LLMs, while achieving comparable performance with extremely great memory reduction and GPU-hour reduction. Subsequently, ZO-AdaMU (Jiang et al., 2023) improved ZO-SGD and adapted the simulated perturbation with momentum in its stochastic approximation method. Both of these two optimizers provide researchers with a new and promising technique for fine-tuning large models.
@@ Appendix B Detailed Convergence Analysis
By the update rule of $\theta_{t}$ and Assumption 4.1, we have
$\mathbb{E}\left[\mathcal{L}(\theta_{t+1})\mid\theta_{t}\right]$
$\leq$ $\mathcal{L}(\theta_{t})-\eta_{t}\mathbb{E}\left[\langle\nabla\mathcal{L}(\theta_{t}),g_{\mu}(\theta_{t})\rangle\right]+\frac{L\eta_{t}^{2}}{2}\mathbb{E}\left[\|g_{\mu}(\theta_{t})\|^{2}\right]$
$\leq$ $\mathcal{L}(\theta_{t})-\eta_{t}\|\nabla\mathcal{L}(\theta_{t})\|_{\Sigma_{t}}^{2}+\eta_{t}\mathcal{O}\left(\mu\|\nabla\mathcal{L}(\theta_{t})\|\right)$
$+2\eta_{t}^{2}L\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\|\nabla\mathcal{L}(\theta_{t})\|_{\Sigma_{t}}^{2}$
@@
$\leq$ $\mathcal{L}(\theta_{t})-\frac{\eta_{t}}{2}\|\nabla\mathcal{L}(\theta_{t})\|_{\Sigma_{t}}^{2}+2\eta_{t}^{2}L\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\|\nabla\mathcal{L}(\theta_{t})\|_{\Sigma_{t}}^{2}$
@@
$=$ $\mathcal{L}(\theta_{t})-\frac{\eta_{t}}{2}\left(1-4\eta_{t}L(\mathrm{tr}(\Sigma_{t})+\beta_{u})\right)\|\nabla\mathcal{L}(\theta_{t})\|_{\Sigma_{t}}^{2}$
@@
$\leq$ $\mathcal{L}(\theta_{t})-\frac{\eta_{t}}{4}\|\nabla\mathcal{L}(\theta_{t})\|_{\Sigma_{t}}^{2}+2\eta_{t}^{2}L\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\sigma^{2}+\mathcal{O}(\mu^{2}),$
@@
$\mathbb{E}\left[\sum_{t=1}^{T}\frac{\eta_{t}}{4}\|\nabla\mathcal{L}(\theta_{t})\|_{\Sigma_{t}}^{2}\right]\leq\sum_{t=1}^{T}\left(\mathcal{L}(\theta_{t})-\mathcal{L}(\theta_{t+1})\right)$
@@
$=$ $\mathcal{L}(\theta_{1})-\mathcal{L}(\theta_{T+1})+2\eta_{t}^{2}L\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\sigma^{2}+\mathcal{O}(T\mu^{2})$
$\leq$ $\mathcal{L}(\theta_{1})-\mathcal{L}(\theta_{*})+2\eta_{t}^{2}L\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\sigma^{2}+\mathcal{O}(T\mu^{2}).$
@@
$\mathbb{E}\left[\|\nabla\mathcal{L}(\theta_{\mbox{out}})\|^{2}\right]=\frac{1}{T}\sum_{t=1}^{T}\|\nabla\mathcal{L}(\theta_{t})\|^{2}\leq\frac{1}{T\beta_{\ell}}\sum_{t=1}^{T}\|\nabla\mathcal{L}(\theta_{t})\|_{\Sigma_{t}}^{2}$
$\leq$ $\frac{4(\mathcal{L}(\theta_{1})-\mathcal{L}(\theta_{*}))}{T\beta_{\ell}\eta}+\frac{8\eta L\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)}{T\beta_{\ell}}\sigma^{2}+\mathcal{O}(\mu^{2})$
$=$ $\frac{32L\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)(\mathcal{L}(\theta_{1})-\mathcal{L}(\theta_{*}))}{\sqrt{T}\beta_{\ell}}+\frac{\sigma^{2}}{T^{3/2}\beta_{\ell}}+\mathcal{O}\left(\mu^{2}\right),$
@@
Eq. (11) shows that once we choose the step size $\eta$ properly, $\mathcal{L}(\theta_{t+1})$ will be less than $\mathcal{L}(\theta_{t})$ in expectation up to some noises of order $\mu^{2}$. Specifically, if set $\eta=\frac{1}{8\sqrt{T}L(\max_{t}\mathrm{tr}(\Sigma_{t})+\beta_{u})}$, Eq. (12) implies that we can find an solution $\theta_{\mbox{out}}$ such that $\mathbb{E}\left[\|\nabla\mathcal{L}(\theta_{\mbox{out}})\|^{2}\right]\leq\epsilon^{2}$ in $\mathcal{O}(\epsilon^{-4})$ iterations. This rate matches the one of (Ghadimi & Lan, 2013).
@@
We assume that Assumption 4.2 and Assumption 4.3 hold. Then, $g_{\mu}(\theta_{t})$ defined in Eq. (10) has the following properties:
@@
$=$ $\sum_{i=1}^{b}\frac{\mathcal{L}(\theta_{t}+\mu\Sigma_{t}^{1/2}u_{i},\mathcal{B}_{t})-\mathcal{L}(\theta_{t}-\mu\Sigma_{t}^{1/2}u_{i},\mathcal{B}_{t})}{2b\mu}\Sigma_{t}^{1/2}u_{i}$
$=$ $\sum_{i=1}^{b}\frac{2\mu\nabla^{\top}\mathcal{L}(\theta_{t},\mathcal{B}_{t})\Sigma_{t}^{1/2}u_{i}+\mathcal{O}(\mu^{2})}{2b\mu}\Sigma_{t}^{1/2}u_{i}$
$=$ $\frac{1}{b}\sum_{i=1}^{b}\Sigma_{t}^{1/2}u_{i}u_{i}^{\top}\Sigma_{t}^{1/2}\nabla\mathcal{L}(\theta_{t},\mathcal{B}_{t})+\mathcal{O}(\mu).$
@@
$\mathbb{E}\left[g_{\mu}(\theta_{t})\right]=\Sigma_{t}\nabla\mathcal{L}(\theta_{t})+\mathcal{O}(\mu).$ (13)

### v3

@@ 4.3 Training with Non-Differentiable Objectives
Our proposed HiZOO employs gradient estimation to update parameters, allowing for the use of non-differentiable objectives for training. Following the setting of MeZO [32], we conduct extensive experiments using F1 as optimization objective. The results presented in Table 4 indicate that our method outperforms MeZO by 6.36% on F1 on average.
Table 3: Experiments on OPT-30B (we use FT and prefix-tuning, report the best of them) and OPT-66B (we use prefix-tuning).
Task SST-2 RTE WSC WIC Average
30B MeZO 90.6 66.4 63.5 59.1 69.9
30B HiZOO 91.2 69.3 63.5 60.2 71.0
30B HiZOO-L 91.1 68.9 63.5 59.8 70.8
66B MeZO 93.6 66.4 57.7 58.6 69.0
66B HiZOO 93.6 71.5 60.6 61.1 71.7
66B HiZOO-L 93.6 71.0 60.3 60.9 71.4
Table 4: Experiments on non-differentiable optimization objectives (F1). For classification ($k=512$), we use full-parameter tuning and for SQuAD (1,000 examples), we use prefix tuning.
@@
4.4 Memory Usage and Time Efficiency Analysis
Figure 5: GPU memory consumption with different OPT models and tuning methods on MultiRC (400 tokens per example on average). More details can be found in Appendix F.
[…]
Time Efficiency    In Appendix F, we analyse the wall-clock time efficiencies and show that HiZOO spends 1.5×1.5$\times$1.5 × time per step compared with MeZO, mainly from the extra forward pass. However, HiZOO reduces total number of forward passes required for convergence. For example, HiZOO achieves a 8×8$\times$8 × and 4×4$\times$4 × speedup on SST2 and MNLI tasks.
4.5 Comparison with other ZO variants
We also compare our HiZOO with a broader array of ZO optimization techniques [59]. As shown in Table 5, our HiZOO outperforms all other ZO methods. Compared with ZO-Adam who leverages second-order moment to guide gradient descent, our HiZOO-L achieves a notable 4.3% absolute improvement, while using 50% of the GPU memory.
Table 5: Performance comparison on SST2(Robert-Large and OPT-1.3B) and COPA(OPT-13B) using different ZO methods. Memory and runtime cost are multiples of ZO-SGD.
Model/Task Roberts-Large OPT-1.3B OPT-13B Average Memory Runtime
FT prefix FT prefix FT prefix
ZO-SGD 89.4 90.0 90.8 91.4 90.0 79.0 88.4 1.0x 1.0x
ZO-SGD-MMT 89.6 89.1 85.2 91.2 87.0 85.0 87.8 1.56x 1.0x
ZO-SGD-Cons 89.6 89.1 88.3 88.1 82.0 84.0 86.8 1.0x 2.49x
ZO-SGD-Sign 52.5 53.6 87.2 89.5 80.0 78.0 73.4 1.0x 1.0x
ZO-Adam 89.8 90.2 84.4 91.4 82.0 79.0 86.1 2.47x 1.04x
HiZOO 93.2 92.7 90.7 91.4 88.0 87.0 90.5 2.04x 1.37x
HiZOO-L 92.5 92.7 90.7 91.4 88.0 87.0 90.4 1.12x 1.39x
4.6 Hyperparameter Analysis
@@
Influence of Smooth Scale $\alpha_{t}$ in EMA    To assess the robustness of the optimizer, a grid search is conducted to evaluate the sensitivity of the hyper-parameter $\alpha_{t}$ on RoBERTa-large (350M). Figure 7 illustrates that the smoothing scale $\alpha_{t}$ is instrumental in controlling the convergence rate during training. As $\alpha_{t}$ is incrementally increased from zero, the training loss decreases faster. However, too large $\alpha_{t}$ values may impede convergence or even cause training to fail due to gradient explosion.
Influence of Estimation Times $n$ Per Step    We also propose a variant of HiZOO in Appendix D.1: HiZOO-multi, which has $n\textgreater 1$ per step. As shown in Figure 7, different $n$ maybe doesn’t affect the final accuracy. However, the larger $n$ will estimate the diagonal Hessian more accurate per step and accelerate model convergence, reducing the overall training steps. But it will also increase the computation per step. Balancing these factors is crucial for efficient training.
Figure 6: Experiments on influence of EMA $\alpha_{t}$ for hessian in Eq. (6). We use HiZOO (prefix) to fine-tune Roberta-large on SNLI. More results can be found in Appendix H.1.
@@
Figure 7: Loss curves on Function (a) using the variant HiZOO-multi with different estimation times $n$ per step. Trajectory visualization can be found in Appendix D.1.
@@ 5 Conclusion
In this work, we introduce HiZOO, which is the first ZOO that incorporates diagonal Hessian for fine-tuning LLMs. By introducing one more forward pass, HiZOO can handle heterogeneous curvatures across different parameter dimensions. We provide theoretical analysis and visualize the optimization trajectories to explore how it works. Further experiments on LLMs show that HiZOO converges in much fewer steps than MeZO and achieves better performance, across various models (RoBERTa-large, OPT, Llama3, Phi-2) and methods (FT, LoRA and prefix). In addition, HiZOO increases extra memory cost because of the storage of Hessian. We explore a memory efficient implementation (HiZOO-L) and there are many other methods worth trying, such as quantization [10], FlashAttention [8].
@@ A.1 First-order Optimizer Used in LLMs
Optimization methods have consistently been a popular research domain, encompassing techniques such as Gradient Descent (GD), Momentum, Adagrad [12], ADADELTA [54], and Newton’s method, which have been instrumental in advancing fields like computer vision. However, the emergence of large-scale models, characterized by their massive parameter counts and intricate architectures, has challenged the efficacy of conventional optimization methods for training tasks. Amidst this landscape, Adam [22] has emerged as the preferred choice for its ability to rapidly converge, making it particularly suitable for the training and fine-tuning large models. Then AdamW [29] was proposed to add a weight decay coefficient to alleviate over-fitting. Notwithstanding these advancements, a limitation persists with these optimizers: they have an implicit batch size ceiling. Exceeding this threshold can provoke extreme gradient updates, thus impeding the convergence rate of the models. This bottleneck is particularly problematic in the context of large-model training, which typically necessitates substantial batch sizes. To circumvent this constraint, LAMB [53] was devised to apply principled layer-wise adaptation strategy to accelerate the training of large models employing large batches.
@@ A.2 Hessian Based First-Order Optimizer
Compared with first-order optimizers, second-order optimizer considers second-order information in the process of gradient calculation. As a result, it has more abundant information to guide gradient descent and is considered to be more promising. Previous studies utilized curvature information to pre-condition the gradient [4, 34, 7]. Subsequently, Magoulas et al. [31] applied diagonal Hessian as the pre-conditioner, which greatly promotes the landing of second-order optimizer in the field of deep learning. Martens [33] approximated the Hessian with conjugate gradient. Schaul et al. [40] utilized diagonal Hessian to automatically adjust the learning rate of SGD during training. Another work [36] extended natural gradient descent to incorporate second order information alongside the manifold information and used a truncated Newton approach for inverting the metric matrix instead of using a diagonal approximation of it. EVA  [56] proposed to use the Kronecker factorization of two small vectors to approximated the Hessian, which significantly reduces memory consumption. AdaHessian [50] incorporates an approximate Hessian diagonal, with spatial averaging and momentum to precondition the gradient vector.
Although great progress has been made in the research of second-order optimizer, it has not been widely used because of the extra computation and memory cost when gradient updating, and this situation is extremely serious in the training of large language models. Based on the above dilemma, recent works [2, 14] proposed to offload Hessian computation to CPUs and utilized ResNets and very large batch size to approximate the Fisher information matrix. Sophia [26] was the first to apply second-order optimizer and achieve a speed-up on large language models in total compute successfully.
@@ A.3 Zeroth-Order Optimizer
Zeroth-order optimization, is also known as derivative-free or black-box optimization. There have been many one-point gradient estimators in past works [13, 43, 46, 42, 20, 1, 38, 47]. However, cursory experiments with one such promising estimator [58] reveal that SPSA outperforms other methods.
In previous works, it appears in a wide range of applications where either the objective function is implicit or its gradient is impossible or too expensive to compute. For example, methods [44, 18] consider derivative-free distributed algorithms for non-convex multi-agent optimization. ZO-BCD[5], ZOO[6], ZO-signSGD [27] and ZO-HessAware [52] utilize zeroth-order stochastic optimization to generate black-box adversarial example in deep learning.
Beyond that, MeZO [32] firstly adapted the classical ZO-SGD method to fine-tune LLMs, while achieving comparable performance with extremely great memory reduction and GPU-hour reduction. Subsequently, ZO-AdaMU [21] improved ZO-SGD and adapted the simulated perturbation with momentum in its stochastic approximation method. Both of these two optimizers provide researchers with a new and promising technique for fine-tuning large models.
@@ Appendix B Detailed Convergence Analysis
Firstly, our convergence analysis requires the following assumptions:
Assumption B.1.
The objective function $\mathcal{L}(\theta;\mathcal{B})$ is $L$-smooth, which means that for any $\theta_{1},\theta_{2}\in\mathbb{R}^{d}$, it holds that:
$\mathcal{L}(\theta_{2};\mathcal{B})\leq\mathcal{L}(\theta_{1};\mathcal{B})+\langle\nabla\mathcal{L}(\theta_{1};\mathcal{B}),\theta_{2}-\theta_{1}\rangle+\frac{L}{2}\|\theta_{2}-\theta_{1}\|^{2}.$ (10)
Assumption B.2.
The stochastic gradient $\nabla\mathcal{L}(\theta;\mathcal{B})$ has $\sigma^{2}$ variance, which means:
$\mathbb{E}\left[\|\nabla\mathcal{L}(\theta;\mathcal{B})-\nabla\mathcal{L}(\theta)\|^{2}\right]\leq\sigma^{2}.$ (11)
Assumption B.3.
Then we will give the detailed proof for convergence.
@@
By the update rule of $\theta_{t}$ and Assumption B.1, we have
$\mathbb{E}\left[\mathcal{L}(\theta_{t+1};\mathcal{B}_{t+1})\mid\theta_{t}\right]$
$\leq$ $\mathcal{L}(\theta_{t};\mathcal{B}_{t})-\eta_{t}\mathbb{E}\left[\langle\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t}),g_{\mu}(\theta_{t})\rangle\right]+\frac{L\eta_{t}^{2}}{2}\mathbb{E}\left[\|g_{\mu}(\theta_{t})\|^{2}\right]$
$\leq$ $\mathcal{L}(\theta_{t};\mathcal{B}_{t})-\eta_{t}\|\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})\|_{\Sigma_{t}}^{2}+\eta_{t}\mathcal{O}\left(\mu\|\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})\|\right)$
$+2\eta_{t}^{2}L\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\|\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})\|_{\Sigma_{t}}^{2}$
@@
$\leq$ $\mathcal{L}(\theta_{t};\mathcal{B}_{t})-\frac{\eta_{t}}{2}\|\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})\|_{\Sigma_{t}}^{2}+2\eta_{t}^{2}L\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\|\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})\|_{\Sigma_{t}}^{2}$
@@
$=$ $\mathcal{L}(\theta_{t};\mathcal{B}_{t})-\frac{\eta_{t}}{2}\left(1-4\eta_{t}L(\mathrm{tr}(\Sigma_{t})+\beta_{u})\right)\|\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})\|_{\Sigma_{t}}^{2}$
@@
$\leq$ $\mathcal{L}(\theta_{t};\mathcal{B}_{t})-\frac{\eta_{t}}{4}\|\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})\|_{\Sigma_{t}}^{2}+2\eta_{t}^{2}L\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\sigma^{2}+\mathcal{O}(\mu^{2}),$
@@
$\mathbb{E}\left[\sum_{t=1}^{T}\frac{\eta_{t}}{4}\|\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})\|_{\Sigma_{t}}^{2}\right]\leq\sum_{t=1}^{T}\left(\mathcal{L}(\theta_{t};\mathcal{B}_{t})-\mathcal{L}(\theta_{t+1};\mathcal{B}_{t+1})\right)$
@@
$=$ $\mathcal{L}(\theta_{1};\mathcal{B}_{1})-\mathcal{L}(\theta_{T+1};\mathcal{B}_{T+1})+2\eta_{t}^{2}L\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\sigma^{2}+\mathcal{O}(T\mu^{2})$
$\leq$ $\mathcal{L}(\theta_{1};\mathcal{B}_{1})-\mathcal{L}(\theta_{*};\mathcal{B}_{*})+2\eta_{t}^{2}L\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\sigma^{2}+\mathcal{O}(T\mu^{2}).$
@@
$\mathbb{E}\left[\|\nabla\mathcal{L}(\theta_{\mbox{out}};\mathcal{B})\|^{2}\right]=\frac{1}{T}\sum_{t=1}^{T}\|\nabla\mathcal{L}(\theta_{t};\mathcal{B})\|^{2}\leq\frac{1}{T\beta_{\ell}}\sum_{t=1}^{T}\|\nabla\mathcal{L}(\theta_{t};\mathcal{B})\|_{\Sigma_{t}}^{2}$
$\leq$ $\frac{4(\mathcal{L}(\theta_{1};\mathcal{B})-\mathcal{L}(\theta_{*};\mathcal{B}))}{T\beta_{\ell}\eta}+\frac{8\eta L\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)}{T\beta_{\ell}}\sigma^{2}+\mathcal{O}(\mu^{2})$
$=$ $\frac{32L\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)(\mathcal{L}(\theta_{1};\mathcal{B})-\mathcal{L}(\theta_{*};\mathcal{B}))}{\sqrt{T}\beta_{\ell}}+\frac{\sigma^{2}}{T^{3/2}\beta_{\ell}}+\mathcal{O}\left(\mu^{2}\right),$
@@
Eq. (8) shows that once we choose the step size $\eta$ properly, $\mathcal{L}(\theta_{t+1};\mathcal{B}_{t+1})$ will be less than $\mathcal{L}(\theta_{t};\mathcal{B}_{t})$ in expectation up to some noises of order $\mu^{2}$. Specifically, if set $\eta=\frac{1}{8\sqrt{T}L(\max_{t}\mathrm{tr}(\Sigma_{t})+\beta_{u})}$, Eq. (9) implies that we can find an solution $\theta_{\mbox{out}}$ such that $\mathbb{E}\left[\|\nabla\mathcal{L}(\theta_{\mbox{out}};\mathcal{B})\|^{2}\right]\leq\epsilon^{2}$ in $\mathcal{O}(\epsilon^{-4})$ iterations. This rate matches the one of [15].
@@
We assume that Assumption B.2 and Assumption B.3 hold. Then, $g_{\mu}(\theta_{t})$ defined in Eq. (7) has the following properties:
@@
$=$ $\sum_{i=1}^{b}\frac{\mathcal{L}(\theta_{t}+\mu\Sigma_{t}^{1/2}u_{i};\mathcal{B}_{t})-\mathcal{L}(\theta_{t}-\mu\Sigma_{t}^{1/2}u_{i};\mathcal{B}_{t})}{2b\mu}\Sigma_{t}^{1/2}u_{i}$
$=$ $\sum_{i=1}^{b}\frac{2\mu\nabla^{\top}\mathcal{L}(\theta_{t};\mathcal{B}_{t})\Sigma_{t}^{1/2}u_{i}+\mathcal{O}(\mu^{2})}{2b\mu}\Sigma_{t}^{1/2}u_{i}$
$=$ $\frac{1}{b}\sum_{i=1}^{b}\Sigma_{t}^{1/2}u_{i}u_{i}^{\top}\Sigma_{t}^{1/2}\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})+\mathcal{O}(\mu).$
@@
$\mathbb{E}\left[g_{\mu}(\theta_{t})\right]=\Sigma_{t}\nabla\mathcal{L}(\theta_{t};\mathcal{B}_{t})+\mathcal{O}(\mu).$ (12)